# Memory-Augmented Quantum Reservoir Computing

### v1

@@ III.4 Simulations with Rydberg atoms
where $\Omega_{j}(t)$ is the Rabi frequency, $\varphi_{j}(t)$ is the laser phase, $\Delta_{j}(t)$ is the detuning and $V_{jk}$ is the interaction term between atoms $j$ and $k$. The term $V_{jk}=\frac{C_{6}}{|\mathbf{x}_{j}-\mathbf{x}_{k}|^{6}}$ describes the Rydberg interaction, specifically the van der Waals interaction, between atoms $j$ and $k$. Here, $\mathbf{x}_{j}$ denotes the position of the $j$-th atom, and the denominator $|\mathbf{x}_{j}-\mathbf{x}_{k}|$ represents the distance between atoms $j$ and $k$. The coefficient $C_{6}$ is a constant that characterizes the strength of the interaction and depends on the specific Rydberg states being used. For our simulation, the default value of $C_{6}$ is set to $862690\times 2\pi\ \text{MHz}\ \mu\text{m}^{6}$. This value corresponds to the interaction strength for the Rydberg state $|r\rangle=|70S_{1/2}\rangle$ of ${}^{87}\text{Rb}$ atoms. According to [39], we set $\Omega=2\pi\times 4.2\ \text{MHz}$.
Both the phases $\varphi_{j}$ and the parameters $\Delta_{j}$ are, in principle, capable of encoding the input information. We assesed the performance of our reservoir computing algorithm using both these encoding methods, and the set of observables $\vec{m}^{A}$. The information about the $k$-th value of the input sequence $\{s_{k}\}$ is encoded homogenously first in the value of the detuning parameter ${\Delta_{j}}_{k}={\Delta}_{k}$, by fixing the value of the phases to $\varphi^{*}$ (Fig. 7), and subsequently in the phase ${\varphi_{j}}_{k}={\varphi_{k}}$, by fixing the value of ${\Delta_{j}}_{k}$ to ${\Delta^{*}}$ (Fig. 8). In the first case, it is obtained that the maximum value of the $VPT_{Mean}$ occurs when the influence of the spin component along the $x$-axis is nullified, and the component along the $y$-axis is maximized, thereby optimizing the interaction of the system with the external electromagnetic field. Suboptimal values are obtained for phases near $43\pi/$ 4. In the second case, the dependence on the hyperparameter $\Delta^{*}$ is more critical, the best performance is peaked at $\Delta^{*}=5\Omega$ for a specific evolution time $\Delta t=2$.
@@ IV Conclusions
We demonstrated the efficiency of our model by applying it to two distinct physical systems: a fully connected Ising model in a transverse field and a Rydberg atom array. Both platforms showcased the versatility of our method, as they were able to successfully process information in a variety of benchmark tasks, including linear memory, NARMA, and the prediction of the chaotic Mackey-Glass time series. In particular, the Valid Prediction Time (VPT) obtained from our model outperformed several standard approaches reported in the literature, demonstrating the robustness of our quantum-classical hybrid method.
The key innovation of our approach lies in the introduction of two operators in the classical post-processing phase, allowing us to extend arbitrarily the classical system’s memory while maintaining the benefits of quantum information processing. This hybrid approach provides a promising pathway for future reservoir computing applications, particularly in cases where quantum resources are limited or expensive to replicate experimentally.
@@ V Acknowledgements
JS acknowledges the contribution from PRIN (Progetti di Rilevante Interesse Nazionale) TURBIMECS - Turbulence in Mediterranean cyclonic events, grant n. 2022S3RSCT CUP H53D23001630006, CUP Master B53D23007500006.

### v2

@@ III.4 Simulations with Rydberg atoms
where $\Omega_{j}(t)$ is the Rabi frequency, $\varphi_{j}(t)$ is the laser phase, $\Delta_{j}(t)$ is the detuning and $V_{jk}$ is the interaction term between atoms $j$ and $k$. The term $V_{jk}=\frac{C_{6}}{|\mathbf{x}_{j}-\mathbf{x}_{k}|^{6}}$ describes the van der Waals interaction, between atoms $j$ and $k$. Here, $\mathbf{x}_{j}$ denotes the position of the $j$-th atom, and the denominator $|\mathbf{x}_{j}-\mathbf{x}_{k}|$ represents the distance between atoms $j$ and $k$. The atoms are placed equally spaced with a minimum distance set to $7\mu\text{m}$ [49]. The coefficient $C_{6}$ is a constant that characterizes the strength of the interaction and depends on the specific Rydberg states being used. For our simulation, the default value of $C_{6}$ is set to $862690\times 2\pi\ \text{MHz}\ \mu\text{m}^{6}$. This value corresponds to the interaction strength for the Rydberg state $|r\rangle=|70S_{1/2}\rangle$ of ${}^{87}\text{Rb}$ atoms. According to Ref. [49], we further set $\Omega=2\pi\times 4.2\ \text{MHz}$.
Both the phases $\varphi_{j}$ and the parameters $\Delta_{j}$ are, in principle, capable of encoding the input information. We assessed the performance of our reservoir computing algorithm using both of these encoding methods, and the set of observables $\vec{m}^{A}$. The information about the $k$-th value of the input sequence $\{s_{k}\}$ is encoded homogeneously, first in the value of the detuning parameter ${\Delta_{j}}_{k}={\Delta}_{k}$, by fixing the value of the phases to $\varphi^{*}$ (Fig. 7), and subsequently in the phase ${\varphi_{j}}_{k}={\varphi_{k}}$, by fixing the value of ${\Delta_{j}}_{k}$ to ${\Delta^{*}}$ (Fig. 8). In the first case, it is obtained that the maximum value of the $VPT_{Mean}$ occurs when the influence of the spin component along the $x$-axis is nullified, and the component along the $y$-axis is maximized, thereby optimizing the interaction of the system with the external electromagnetic field. Suboptimal values are obtained for phases near $43\pi/$ 4. In the second case, the dependence on the hyper-parameter $\Delta^{*}$ is more critical, the best performance is peaked at $\Delta^{*}=5\Omega$ for a specific evolution time $\Delta t=2$.
@@ IV Conclusions
We demonstrated the efficiency of our model by applying it to two distinct physical systems: a fully connected Ising model in a transverse field and a Rydberg atom array. Both platforms showcased the versatility of our method, as they were able to successfully process information in a variety of benchmark tasks, including linear memory, NARMA, and the prediction of the chaotic Mackey-Glass time series. In particular, the Valid Prediction Time (VPT) obtained from our model outperformed several standard approaches reported in the literature, demonstrating the advanced predictive capabilities of our quantum-classical hybrid method.
The key innovation of our approach lies in the introduction of two operators in the classical post-processing phase, allowing us to extend arbitrarily the classical system’s memory while maintaining the benefits of quantum information processing. The proposed method allows a continuous transition from an Extreme Learning Machine when $\gamma=0$ to QRC with a tunable memory via the parameter $\gamma$. This tunability enables optimizing the system’s memory capacity according to the specific requirements of the target applications. This hybrid approach provides a promising pathway for future reservoir computing applications, particularly in cases where quantum resources are limited or expensive to replicate experimentally.
@@ V Acknowledgements
JS acknowledges the contribution from PRIN (Progetti di Rilevante Interesse Nazionale) TURBIMECS - Turbulence in Mediterranean cyclonic events, grant n. 2022S3RSCT CUP H53D23001630006, CUP Master B53D23007500006, CM from ICSC – Italian Research Center on High Performance Computing, Big Data and Quantum Computing, funded by European Union – NextGenerationEU, PUN: B93C22000620006.